# Bogomolov property and Galois representations

### v3

@@ 1. Introduction
Let $\alpha$ be an algebraic number of degree $d$. We denote by $h(\alpha)$ the absolute and logarithmic Weil height. Assuming that $\alpha$ is not a root of unity, D. H. Lehmer suggested that $dh(\alpha)$ can be bounded below uniformly in $\alpha$. Known as “Lehmer’s problem”, this question is still open. Nevertheless, in some special cases, not only Lehmer’s conjecture is true, but it can also be sharpened. This different point of view was considered for the first time in [5], where the authors introduced the notion of fields with Bogomolov property, (property (B) for short), where the height is uniformly bounded below outside torsion points. By Northcott’s Theorem, number fields are trivial examples of fields with (B). The compositum ${\mathbb{Q}}^{\rm tr}$ of all totally real fields is a less trivial example (apply [24, Corollary $1^{\prime}$] to the linear polynomial $P(z)=z-\alpha$ to get the optimal lower bound). In [5] the authors also show that fields with bounded local degrees at some finite place has Property (B). Note that by a result of [9] a Galois extension has uniformly bounded local degrees at every finite place if and only if its Galois group has finite exponent. Thus (infinite) Galois extensions with Galois group of bounded exponent have Property (B).
@@
Let $\Omega$ be a separated topological group, ${\mathbb{F}}$ be a subfield of $\overline{\mathbb{Q}}$ and $G_{\mathbb{F}}=\mathrm{Gal}(\overline{\mathbb{Q}}/{\mathbb{F}})$. We say that a continuous homomorphism
$\rho:G_{\mathbb{F}}\longrightarrow\Omega,$
has property (B) if $\mathbb{F}(\rho):=\overline{\mathbb{Q}}^{\ker(\rho)}$ does.
@@
In this vein, [1, Corollary 1.7] has the following consequence. Let ${\mathbb{K}}$ be a number field, $\rho:G_{\mathbb{K}}\to\mathrm{GL}_{n}({\mathbb{F}})$ be an absolutely irreducible representation, and ${\mathbb{L}}=\mathbb{K}(\rho)$. Schur’s Lemma asserts that every intertwining automorphism of $\rho$ is a scalar operator. In particular the image of the center of $G=\mathrm{Gal}({\mathbb{L}}/{\mathbb{K}})$ is contained in the subgroup of scalar matrices of $\mathrm{GL}_{n}({\mathbb{F}})$, so that the reduction map $\tilde{\rho}:G\to\mathrm{PGL}_{n}({\mathbb{F}})$ factors through $G/Z(G)$. Thus, if the image of $\tilde{\rho}$ has finite exponent, the same is true for $G/Z(G)$ and [1, Corollary 1.7] implies that $\rho$ has property (B).
@@
If $f$ is a CM modular form, then $\mathbb{Q}(\rho_{f})$ is an abelian extension of a quadratic field, (see for instance [20, Proposition 4.4]). Thus $\rho_{f}$ has property (B) by the special case $D=1$ of [3, Theorem 1.1].
@@
Our condition $(P2)$ generalizes Habegger’s assumption on the surjectivity of the Galois representation $G_{\mathbb{Q}}\to{\rm Aut}{\mathcal{E}}[p]$. It has been extensively investigated in the literature (see [22, 23, 4, 14, 18]); we summarize here the most relevant results for the purposes of this work.
@@
For a modular form such that $\mathbb{K}_{f}\not=\mathbb{Q}$ the situation concerning the image of $\rho_{f,p}$ is complicated by the possible of presence of inner twists; i.e. the elements of the subgroup $\Gamma\subseteq Aut(\mathbb{K}_{f})$ consisting of those automorphisms $\sigma$ for which $f^{\sigma}$ is equal to the twist of $f$ by some Dirichlet character. Let $\mathbb{K}_{1}$ be the subfield of $\mathbb{K}$ fixed by $\Gamma$. It is proved in [22, Theorem 3.1] that there is an abelian number field $E$ such that for $p$ large enough relatively to $f$, the restriction $\rho_{f,p}|_{G_{E}}$ has image $\widehat{G}(\mathcal{O}_{\mathbb{K}_{1}}\otimes\mathbb{Z}_{p})$.
Therefore, all we can say is that condition $(P2)$ holds for $p$ large enough when $f$ has no inner twists. In the case of trivial nebentypus (the one we are dealing with) it is easy to see that inner twists must arise from a quadratic Dirichlet character; and Ribet proved that they do not appear when $N$ is square free [21, 3.9].
@@
Some other examples in weight $k>2$ with $a_{p}=0$, can be produced when $\mathbb{K}=\mathbb{Q}$. In this case in fact conditions (i),(ii), (iii) in [22, page 192] are automatically fulfilled, and if $a_{p}=0$, then condition (iv) is also satisfied, because the residual representation $\bar{\rho}_{v}$ is irreducible, for every $v|p$. There are several examples of this type with $k=4$, for example
@@ 2. Notations and preliminary results
For a field $k$, $\overline{k}$ denotes an algebraic closure of $k$, and $G_{k}=\mathrm{Gal}(\overline{k}/k)$ the absolute Galois group of $k$, endowed with the profinite topology.
Let $\Omega$ be a separated topological group and $\rho:G_{\mathbb{F}}\to\Omega$ be a continuous homomorphism; we shall denote by $k(\rho)=\overline{k}^{\ker\rho}$ the field cut out by $\rho$. Then $k(\rho)/k$ is a Galois extension whose Galois group is isomorphic to ${\rm Im}(\rho)$. Moreover, since $\ker{\rho}$ is closed, $G_{k(\rho)}=\ker\rho$.
When $k$ is a local or global field and $v$ is a place of $k$, we shall say that $\rho$ is unramified (resp. totally ramified) at $v$ if $k(\rho)$ is. Note that $\rho$ induces an isomorphism $\mathrm{Gal}(k(\rho)/k)\rightarrow{\rm Im}(\rho)$ which as usual we denote by the same letter $\rho$.
From now on we fix a normalized new eigenform $f=\sum_{n}a_{n}q^{n}\in S_{k}(\Gamma_{0}(N))$ for the Hecke algebra of weight $k\geq 2$; then $k$ is an even integer. We recall the relevant definitions. We denote by ${\mathbb{K}}$ the number field generated by the $a_{n}$’s and by $\mathcal{O}$ its integer ring. A construction due to Shimura and Deligne associates to $f$ a family of compatible Galois representations:
@@
Let ${\mathbb{L}}$ be a (possible infinite) Galois extension. We now make the first assumption on ${\mathbb{L}}$ which allows to globalise the metric inequality at ramified primes. We follow some ideas from [3, Lemma 4.2] and [15, Proof of Proposition 6.1].
@@
(3.1) $\epsilon_{p}(\tau)=g\quad\hbox{with }g\in\mathbb{Z},g>1,$
@@
Let’s come back to our representations $\rho_{v}$, with $v\mid p$. Let as assume $p\nmid N$ (Assumption 1.3 $(P0)$). By [25, Theorem 1.2.4 (ii)], ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ is crystalline and the characteristic polynomial of $\varphi$ on the associated $\varphi$-filtered module is equal to $X^{2}-a_{p}X+p^{k-1}$. Moreover, representations associate to modular forms of weight $k$ have Hodge-Tate weight $(0,k-1)$, (see [11]). It follows that the filtered module associated to ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ is $D_{k^{\prime},a,\chi}$ with $k^{\prime}=k$, $a=a_{p}$, and $\chi(p)=1$ or $-1$. Now assume also $a_{p}=0$ (Assumption 1.3 $(P1)$). Since $V_{k^{\prime},a,\chi}\simeq V_{k^{\prime},-a,\chi\mu_{-1}}$, where $\mu_{-1}$ is the unramified quadratic character of $G_{{\mathbb{Q}}_{p}}$, in both cases $\rho_{v}|_{G_{\mathbb{Q}_{p}}}$ is isomorphic to $V_{k,0}$. Taking into account the explicit description of $V_{k,0}$ in [6, Prop. 3.2] we obtain:
@@
$V_{k,0}\simeq\left(\mathrm{ind}_{G_{\mathbb{Q}_{q}}}^{G_{\mathbb{Q}_{p}}}\epsilon_{2}^{k-1}\right)\otimes\mu_{\sqrt{-1}},$
@@
- $\epsilon_{2}:G_{\mathbb{Q}_{q}}\to{\mathbb{Z}}_{q}^{\times}$ is one of the two continuous characters characterized by the property that their composition with the injection $\mathbb{Q}_{q}^{\times}\hookrightarrow G_{\mathbb{Q}_{q}}^{ab}$ (given by local class field theory) is trivial on $p$ and is one of the two natural embeddings when restricted to $\mathbb{Z}_{q}^{\times}$;
@@
Let $p$ be a rational prime satisfying Assumptions 1.3 $(P0)$, $(P1)$. By Propositions 4.1, for every prime $v$ of $\mathcal{O}$ above $p$, ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ is a twist by an unramified character of the induced representation of a power of a character arising from the Tate module of a Lubin-Tate formal group over $\mathbb{Z}_{q}$ (see [10, Théorème I.2.1]). This is exactly the same situation occuring for the representation of $G_{{\mathbb{Q}}_{p}}$ on the $p$-adic Tate module of an elliptic curve ${\mathcal{E}}$ over ${\mathbb{Q}}$ having $p$ as a supersingular prime, and this is the key point of the fundamental Galois theoretic analysis of the $p$-power torsion points on $E$ in [15, Lemma 3.3]. The proof of this lemma rests on the theory of Lubin-Tate modules and its relation to local class field theory and thus only depends on the local $p$-adic representation. We provide here a reformulation suitable for our situation, without any reference to the theory of elliptic curves.
@@
Let $n\geq 1$. We denote by $\epsilon_{2,n}$ the composition of $\epsilon_{2}$ with the reduction modulo $p^{n}$.
@@
    The extension ${\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}$ is totally ramified and abelian of degree $\delta q^{n-1}$, and
    $\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q})\simeq\{g^{k-1}\,|\,g\in({\mathbb{Z}}_{q}/p^{n}{\mathbb{Z}}_{q})^{\times}\}\simeq{\mathbb{Z}}/\delta{\mathbb{Z}}\times\left({\mathbb{Z}}/p^{n-1}{\mathbb{Z}}\right)^{2}.$
@@
    $G_{i}({\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}))=\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}(\epsilon_{2,j}^{k-1})).$
@@
By [10, Théorème I.2.1], $\epsilon_{2}$ is up to isogeny the Tate module of a Lubin–Tate formal group $\mathcal{F}$ over ${\mathbb{Z}}_{q}$. Then ${\mathbb{Q}}_{q}(\epsilon_{2,n})={\mathbb{Q}}_{q}(\mathcal{F}(n))$, where $\mathcal{F}(n)$ is the group of $p^{n}$-division points of $\mathcal{F}$. Then in the case $k=2$ (i) and (ii) follow from [19, Theorem V.5.4] and [19, Proposition V.6.1].
Assume now $k>2$. The extension ${\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}$ is the subextension of ${\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}$ cut out by the $(k-1)$-th power of $\epsilon_{2,n}$. By assertion (i) in the special case $k=2$, and since $p\nmid k-1$ by Assumption 1.3 $(P3)$, we see that $\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1}))$ is the subgroup of ${\mathbb{Z}}/(q-1){\mathbb{Z}}$ of index $d:=(q-1,k-1)$. Assertion (i) follows.
We now prove (ii). The group $\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1}))$ has order $d$. By Herbrand Theorem [19, Ch. II, Theorem 10.7], $G_{r}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q})$ is mapped onto $G_{\eta(r)}({\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q})$ by the restriction map
$\mathrm{res}\colon\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q})\longrightarrow\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}).$
@@
$\eta(r)=\frac{1}{d}\sum_{\sigma\in\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1}))}\min\{i(\sigma),r+1\}-1$
with $i(\sigma)=v(\sigma(x)-x)$ for a generator $x$ of the ring of integers of ${\mathbb{Q}}_{q}(\epsilon_{2,n})$ over ${\mathbb{Q}}_{q}(\epsilon_{2,n})$. For $\sigma\in\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1}))$ we have
@@
because $G_{1}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q})$ is a $p$-group and $\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1}))$ has order dividing $q-1$ and thus $G_{1}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q})\cap\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1}))$ is trivial. Therefore
@@
$G_{i}({\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}))$ $=\mathrm{res}\left(G_{d\,i}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q})\right)$
$=\mathrm{res}\left({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}(\epsilon_{2,j})\right)\quad\hbox{(by the case $k=2$)}$
$=\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}(\epsilon_{2,j}^{k-1})).$
@@
$\rho_{v|G_{{\mathbb{Q}}_{q}}}\simeq\left(\mathrm{ind}_{G_{{\mathbb{Q}}_{q}}}^{G_{{\mathbb{Q}}_{p}}}\epsilon_{2}^{k-1}\right)\otimes\mu_{\sqrt{-1}}.$
@@
(4.4) $\left(\mathrm{ind}_{G_{{\mathbb{Q}}_{q}}}^{G_{{\mathbb{Q}}_{p}}}\epsilon_{2}^{k-1}\right)_{|G_{{\mathbb{Q}}_{q}}}=\epsilon_{2}^{k-1}\oplus{\epsilon^{\prime}_{2}}^{k-1},$
where $\epsilon^{\prime}_{2}$ is the character obtained from $\epsilon_{2}$ by composing with the algebraic conjugation of ${\mathbb{Q}}_{q}/{\mathbb{Q}}_{p}$. Let for short $\nu=\epsilon_{2}^{k-1}$, $\omega=\mu_{\sqrt{-1}|G_{{\mathbb{Q}}_{q}}}$ and remark that $\omega$ takes values in {±$\{\pm 1\}$ and thus it is stable by the algebraic conjugation of ${\mathbb{Q}}_{q}/{\mathbb{Q}}_{p}$. Notice also that $-1\in{\mathrm{im}}(\nu)=\left(\mathbb{Z}_{q}^{\times}\right)^{k-1}$, because the weight $k$ is even. Then $\rho_{v|G_{{\mathbb{Q}}_{q}}}$ is conjugate to

### v4

@@ 1. Introduction
Let $\alpha$ be an algebraic number of degree $d$. We denote by $h(\alpha)$ the absolute and logarithmic Weil height. Assuming that $\alpha$ is not a root of unity, D. H. Lehmer suggested that $dh(\alpha)$ can be bounded below uniformly in $\alpha$. Known as “Lehmer’s problem”, this question is still open. Nevertheless, in some special cases, not only Lehmer’s conjecture is true, but it can also be sharpened. This different point of view was considered for the first time in [5], where the authors introduced the notion of fields with Bogomolov property, (property (B) for short), where the height is uniformly bounded below outside torsion points. By Northcott’s Theorem, number fields are trivial examples of fields with (B). The compositum ${\mathbb{Q}}^{\rm tr}$ of all totally real fields is a less trivial example (apply [25, Corollary $1^{\prime}$] to the linear polynomial $P(z)=z-\alpha$ to get the optimal lower bound). In [5] the authors also show that fields with bounded local degrees at some finite place has Property (B). Note that by a result of [9] a Galois extension has uniformly bounded local degrees at every finite place if and only if its Galois group has finite exponent. Thus (infinite) Galois extensions with Galois group of bounded exponent have Property (B).
@@
Let $\Omega$ be a separated topological group, $\mathbb{K}$ be a subfield of $\overline{\mathbb{Q}}$ and $G_{\mathbb{K}}=\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{K})$. We say that a continuous homomorphism
$\rho:G_{\mathbb{K}}\longrightarrow\Omega,$
has property (B) if $\mathbb{K}(\rho):=\overline{\mathbb{Q}}^{\ker(\rho)}$ does.
@@
In this vein, [1, Corollary 1.7] has the following consequence. Let ${\mathbb{K}}$ be a subfield of $\overline{\mathbb{Q}}$, $\kappa$ be any field, $\rho:G_{\mathbb{K}}\to\mathrm{GL}_{n}(\kappa)$ be an absolutely irreducible representation, and ${\mathbb{L}}=\mathbb{K}(\rho)$. Schur’s Lemma asserts that every intertwining automorphism of $\rho$ is a scalar operator. In particular the image of the center of $G=\mathrm{Gal}({\mathbb{L}}/{\mathbb{K}})$ is contained in the subgroup of scalar matrices of $\mathrm{GL}_{n}(\kappa)$, so that the reduction map $\tilde{\rho}:G\to\mathrm{PGL}_{n}(\kappa)$ factors through $G/Z(G)$. Thus, if the image of $\tilde{\rho}$ has finite exponent, the same is true for $G/Z(G)$ and [1, Corollary 1.7] implies that $\rho$ has property (B).
@@
If $f$ is a CM modular form, then $\mathbb{Q}(\rho_{f})$ is an abelian extension of a quadratic field, (see for instance [21, Proposition 4.4]). Thus $\rho_{f}$ has property (B) by the special case $D=1$ of [3, Theorem 1.1].
@@
In Section 6 we also prove a more general statement (Corollary 6.2), claiming that property (B)  holds for $\mathbb{F}(\rho_{f})$, where $\mathbb{F}$ is any Galois algebraic field unramified at $p$ and such that a power of Frobenius is central in $\mathrm{Gal}(\mathbb{F}/\mathbb{Q})$.
@@
Our condition $(P2)$ generalizes Habegger’s assumption on the surjectivity of the Galois representation $G_{\mathbb{Q}}\to{\rm Aut}({\mathcal{E}}[p])$. It has been extensively investigated in the literature (see [23, 24, 4, 14, 18]); we summarize here the most relevant results for the purposes of this work.
@@
For a modular form such that $\mathbb{K}_{f}\not=\mathbb{Q}$ the situation concerning the image of $\rho_{f,p}$ is complicated by the possible of presence of inner twists; i.e. the elements of the subgroup $\Gamma\subseteq\mathrm{Aut}(\mathbb{K}_{f})$ consisting of those automorphisms $\sigma$ for which $f^{\sigma}$ is equal to the twist of $f$ by some Dirichlet character. Let $\mathbb{K}_{1}$ be the subfield of $\mathbb{K}$ fixed by $\Gamma$. It is proved in [23, Theorem 3.1] that there is an abelian number field $E$ such that for $p$ large enough relatively to $f$, the restriction $\rho_{f,p}|_{G_{E}}$ has image $\widehat{G}(\mathcal{O}_{\mathbb{K}_{1}}\otimes\mathbb{Z}_{p})$.
Therefore, all we can say is that condition $(P2)$ holds for $p$ large enough when $f$ has no inner twists. In the case of trivial nebentypus (the one we are dealing with) it is easy to see that inner twists must arise from a quadratic Dirichlet character; and Ribet proved that they do not appear when $N$ is square free [22, 3.9].
@@
Some other examples in weight $k>2$ with $a_{p}=0$, can be produced when $\mathbb{K}=\mathbb{Q}$. In this case in fact conditions (i),(ii), (iii) in [23, page 192] are automatically fulfilled, and if $a_{p}=0$, then condition (iv) is also satisfied, because the residual representation $\bar{\rho}_{v}$ is irreducible, for every $v|p$. There are several examples of this type with $k=4$, for example
@@ 2. Notations and preliminary results
(It should be noted in particular that $\mathbb{F}$ denotes a global field rather than a finite one. On the other hand, in accordance with standard notation, the finite field with $q$ elements will be denoted by $\mathbb{F}_{q}$. This should not cause any confusion to the reader.)
For a field $\kappa$, $\overline{\kappa}$ denotes an algebraic closure of $\kappa$, and $G_{\kappa}=\mathrm{Gal}(\overline{\kappa}/\kappa)$ the absolute Galois group of $\kappa$, endowed with the Krull topology.
Let $\Omega$ be a separated topological group and $\rho:G_{\kappa}\to\Omega$ be a continuous homomorphism; we shall denote by $\kappa(\rho)=\overline{\kappa}^{\ker\rho}$ the field cut out by $\rho$. Then $\kappa(\rho)/\kappa$ is a Galois extension whose Galois group is isomorphic to ${\rm Im}(\rho)$. Moreover, since $\ker{\rho}$ is closed, $G_{\kappa(\rho)}=\ker\rho$.
When $\kappa$ is a local or global field and $v$ is a place of $\kappa$, we shall say that $\rho$ is unramified (resp. totally ramified) at $v$ if $\kappa(\rho)$ is. Note that $\rho$ induces an isomorphism $\mathrm{Gal}(\kappa(\rho)/\kappa)\rightarrow{\rm Im}(\rho)$ which as usual we denote by the same letter $\rho$.
From now on we fix a normalized eigenform $f=\sum_{n}a_{n}q^{n}\in S_{k}(\Gamma_{0}(N))$ for the Hecke algebra of weight $k\geq 2$; then $k$ is an even integer. We recall the relevant definitions. We denote by ${\mathbb{K}}$ the number field generated by the $a_{n}$’s and by $\mathcal{O}$ its integer ring. A construction due to Shimura and Deligne associates to $f$ a family of compatible Galois representations:
@@
Let ${\mathbb{L}}$ be a (possible infinite) Galois extension. We now make the first assumption on ${\mathbb{L}}$ which allows to globalise the metric inequality at ramified primes. We follow some ideas from the proofs of [1, Proposition 4.2] and [15, Proposition 6.1].
@@
(3.1) $\varepsilon_{p}(\tau)=g\quad\hbox{with }g\in\mathbb{Z},g>1,$
@@
Let’s come back to our representations $\rho_{v}$, with $v\mid p$. Let as assume $p\nmid N$ (Assumption 1.3 $(P0)$). By [26, Theorem 1.2.4 (ii)], ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ is crystalline and the characteristic polynomial of $\varphi$ on the associated $\varphi$-filtered module is equal to $X^{2}-a_{p}X+p^{k-1}$. Moreover, representations associate to modular forms of weight $k$ have Hodge-Tate weight $(0,k-1)$, (see [11]). It follows that the filtered module associated to ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ is $D_{k^{\prime},a,\chi}$ with $k^{\prime}=k$, $a=a_{p}$, and $\chi(p)=1$ or $-1$. Now assume also $a_{p}=0$ (Assumption 1.3 $(P1)$). Since $V_{k^{\prime},a,\chi}\simeq V_{k^{\prime},-a,\chi\mu_{-1}}$, where $\mu_{-1}$ is the unramified quadratic character of $G_{{\mathbb{Q}}_{p}}$, in both cases $\rho_{v}|_{G_{\mathbb{Q}_{p}}}$ is isomorphic to $V_{k,0}$. Taking into account the explicit description of $V_{k,0}$ in [6, Prop. 3.2] we obtain:
@@
$V_{k,0}\simeq\left(\mathrm{ind}_{G_{\mathbb{Q}_{q}}}^{G_{\mathbb{Q}_{p}}}\varepsilon_{2}^{k-1}\right)\otimes\mu_{\sqrt{-1}},$
@@
- $\varepsilon_{2}:G_{\mathbb{Q}_{q}}\to{\mathbb{Z}}_{q}^{\times}$ is one of the two continuous characters characterized by the property that their composition with the injection $\mathbb{Q}_{q}^{\times}\hookrightarrow G_{\mathbb{Q}_{q}}^{ab}$ (given by local class field theory) is trivial on $p$ and is one of the two natural embeddings when restricted to $\mathbb{Z}_{q}^{\times}$;
@@
Let $p$ be a rational prime satisfying Assumptions 1.3 $(P0)$, $(P1)$. By Proposition 4.1, for every prime $v$ of $\mathcal{O}$ above $p$, ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ is a twist by an unramified character of the induced representation of a power of a character arising from the Tate module of a Lubin-Tate formal group over $\mathbb{Z}_{q}$ (see [10, Théorème I.2.1]). This is exactly the same situation occuring for the representation of $G_{{\mathbb{Q}}_{p}}$ on the $p$-adic Tate module of an elliptic curve ${\mathcal{E}}$ over ${\mathbb{Q}}$ having $p$ as a supersingular prime, and this is the key point of the fundamental Galois theoretic analysis of the $p$-power torsion points on ${\mathcal{E}}$ in [15, Lemma 3.3]. The proof of this lemma rests on the theory of Lubin-Tate modules and its relation to local class field theory and thus only depends on the local $p$-adic representation. We provide here a reformulation suitable for our situation, without any reference to the theory of elliptic curves.
@@
Let $n\geq 1$. We denote by $\varepsilon_{2,n}$ the composition of $\varepsilon_{2}$ with the reduction modulo $p^{n}$.
@@
    The extension ${\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}$ is totally ramified and abelian of degree $\delta q^{n-1}$, and
    $\mathrm{Gal}({\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q})\simeq\{g^{k-1}\,|\,g\in({\mathbb{Z}}_{q}/p^{n}{\mathbb{Z}}_{q})^{\times}\}\simeq{\mathbb{Z}}/\delta{\mathbb{Z}}\times\left({\mathbb{Z}}/p^{n-1}{\mathbb{Z}}\right)^{2}.$
@@
    $G_{i}({\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}))=\mathrm{Gal}({\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}(\varepsilon_{2,j}^{k-1})).$
@@
By [10, Théorème I.2.1], $\varepsilon_{2}$ is up to isogeny the Tate module of a Lubin–Tate formal group $\mathcal{F}$ over ${\mathbb{Z}}_{q}$. Then ${\mathbb{Q}}_{q}(\varepsilon_{2,n})={\mathbb{Q}}_{q}(\mathcal{F}(n))$, where $\mathcal{F}(n)$ is the group of $p^{n}$-division points of $\mathcal{F}$. Then in the case $k=2$ (i) and (ii) follow from [19, Theorem V.5.4] and [19, Proposition V.6.1].
Assume now $k>2$. The extension ${\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}$ is the subextension of ${\mathbb{Q}}_{q}(\varepsilon_{2,n})/{\mathbb{Q}}_{q}$ cut out by the $(k-1)$-th power of $\varepsilon_{2,n}$. By assertion (i) in the special case $k=2$, and since $p\nmid k-1$ by Assumption 1.3 $(P3)$, we see that $\mathrm{Gal}({\mathbb{Q}}_{q}(\varepsilon_{2,n})/{\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1}))$ is the subgroup of ${\mathbb{Z}}/(q-1){\mathbb{Z}}$ of index $d:=(q-1,k-1)$. Assertion (i) follows.
We now prove (ii). The group $\mathrm{Gal}({\mathbb{Q}}_{q}(\varepsilon_{2,n})/{\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1}))$ has order $d$. By Herbrand Theorem [19, Ch. II, Theorem 10.7], the ramification group $G_{r}({\mathbb{Q}}_{q}(\varepsilon_{2,n})/{\mathbb{Q}}_{q})$ is mapped onto $G_{\eta(r)}({\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q})$ by the restriction map
$\mathrm{res}\colon\mathrm{Gal}({\mathbb{Q}}_{q}(\varepsilon_{2,n})/{\mathbb{Q}}_{q})\longrightarrow\mathrm{Gal}({\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}).$
@@
$\eta(r)=\frac{1}{d}\sum_{\sigma\in\mathrm{Gal}({\mathbb{Q}}_{q}(\varepsilon_{2,n})/{\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1}))}\min\{i(\sigma),r+1\}-1$
with $i(\sigma)=v(\sigma(x)-x)$ for a generator $x$ of the ring of integers of ${\mathbb{Q}}_{q}(\varepsilon_{2,n})$ over ${\mathbb{Q}}_{q}(\varepsilon_{2,n})$. For $\sigma\in\mathrm{Gal}({\mathbb{Q}}_{q}(\varepsilon_{2,n})/{\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1}))$ we have
@@
because $G_{1}({\mathbb{Q}}_{q}(\varepsilon_{2,n})/{\mathbb{Q}}_{q})$ is a $p$-group and $\mathrm{Gal}({\mathbb{Q}}_{q}(\varepsilon_{2,n})/{\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1}))$ has order dividing $q-1$ and thus $G_{1}({\mathbb{Q}}_{q}(\varepsilon_{2,n})/{\mathbb{Q}}_{q})\cap\mathrm{Gal}({\mathbb{Q}}_{q}(\varepsilon_{2,n})/{\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1}))$ is trivial. Therefore
@@
$G_{i}({\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}))$ $=\mathrm{res}\left(G_{d\,i}({\mathbb{Q}}_{q}(\varepsilon_{2,n})/{\mathbb{Q}}_{q})\right)$
$=\mathrm{res}\left({\mathbb{Q}}_{q}(\varepsilon_{2,n})/{\mathbb{Q}}_{q}(\varepsilon_{2,j})\right)\quad\hbox{(by the case $k=2$)}$
$=\mathrm{Gal}({\mathbb{Q}}_{q}(\varepsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}(\varepsilon_{2,j}^{k-1})).$
@@
$\rho_{v|G_{{\mathbb{Q}}_{q}}}\simeq\left(\mathrm{ind}_{G_{{\mathbb{Q}}_{q}}}^{G_{{\mathbb{Q}}_{p}}}\varepsilon_{2}^{k-1}\right)\otimes\mu_{\sqrt{-1}}.$
@@
(4.4) $\left(\mathrm{ind}_{G_{{\mathbb{Q}}_{q}}}^{G_{{\mathbb{Q}}_{p}}}\varepsilon_{2}^{k-1}\right)_{|G_{{\mathbb{Q}}_{q}}}=\varepsilon_{2}^{k-1}\oplus{\varepsilon^{\prime}_{2}}^{k-1},$
where $\varepsilon^{\prime}_{2}$ is the character obtained from $\varepsilon_{2}$ by composing with the algebraic conjugation of ${\mathbb{Q}}_{q}/{\mathbb{Q}}_{p}$. Let for short $\nu=\varepsilon_{2}^{k-1}$, $\omega=\mu_{\sqrt{-1}|G_{{\mathbb{Q}}_{q}}}$ and remark that $\omega$ takes values in {±$\{\pm 1\}$ and thus it is stable by the algebraic conjugation of ${\mathbb{Q}}_{q}/{\mathbb{Q}}_{p}$. Notice also that $-1\in{\mathrm{im}}(\nu)=\left(\mathbb{Z}_{q}^{\times}\right)^{k-1}$, because the weight $k$ is even. Then $\rho_{v|G_{{\mathbb{Q}}_{q}}}$ is conjugate to